OpenStudy (anonymous):

Here's the next one! Question: Let f(x) = 9x − 2 and g(x) = −x + 3. Find f(g(x)). Answer Choices: A) -9x - 2 B) -9x + 5 C) -9x + 25 D) -9x + 27

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

Do you want to try it like how you did the last one, then we go over it? :-)

OpenStudy (anonymous):

Yes, I'm pretty I've got this one down. xD

OpenStudy (maddy1251):

Haha, okay, I believe you do, too :)

OpenStudy (anonymous):

Okay, it's either: -9x + 25? My equation was: 9(-x+3)-2

OpenStudy (maddy1251):

Good job. We have f(x) = 9x − 2 and g(x) = −x + 3. We want to find f(g(x)) = f(-x+3)=9(-x+3)-2. =-9x+27-2 -9x+25 is your answer. Excellent excellent!

OpenStudy (anonymous):

Okay lol, I have only 2 more questions left. They look very tricky though. You wouldn't mind helping me with them? Should I close this question or we could discuss it here since before the messaging wasn't all that long.

OpenStudy (maddy1251):

I am willing to help you. Doesn't matter where you post it :-)

OpenStudy (anonymous):

Okay, I'd like it on another post so I just seems more clear to me. Lol, I'll tag you in it. :) Thank you once again!

OpenStudy (maddy1251):

Anytime!! :) I'm glad to be helping :)